The function C(x)=25x + 80000 express the total cost (C)(in $) of manufacturing x units of a product. If the maximum number of units which can be produced equals 20000, state the restricted domain and range for this function.
PLEASE HELP ME FIND SOLUTION TO THAT PROBLEM, THANKS.

Answer by Theo(13342)

c(x) = 25x + 80000.

your x value has to be less than or equal to 20000.
presumably it also has to be greater than 0, because you can't have a negative number of units produced.

so your domain is 0 <= x <= 20,000.

when x is equal to 0, your total cost is 80,000.

when x is equal to 20,000, your total cost is 25 * 20,000 + 80,000 = 580,000.

your range is 80,000 <= y <= 580,000.

the graph of your cost equation is y = 25*x + 80000

you restrictions are that:

x can't be less than 0.
y can't be less than 0.
x can't be greater than 20,000.

based on those restrictions, your graph would look something like this.

$$$

the shaded regions are the value that x and y can't take.

y can't be less than 0.
x can't be less than 0 or greater than 20,000.

the permisible values for your cost function are in the white and on the line of y = 25*x + 80,000.

the limits are between 80,000 and 580,000.